ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 195, 2013

Alarcón, A. and F. J. López Compact complete null curves in complex 3-space

97

Anantharaman, N. and L. Silberman A Haar component for quantum limits on locally symmetric spaces

393

Araujo-Pardo, G., I. Hubard, D. Oliveros and E. Schulte Colorful polytopes and graphs

647

Avilés, A., G. Plebanek and J. Rodríguez Measurability in C(2^{κ}) and Kunen cardinals

1

Banerjee, D. and E. Ghate Adjoint lifts and modular endomorphism algebras

507

Barvinok, A., S. J. Lee and I. Novik Centrally symmetric polytopes with many faces

457

Boumov, S. and V. Drensky Algebraic properties of codimension series of P I-algebras

593

Brandl, R. Conjugacy classes of non-normal subgroups of finite pgroups

473

Brendle, T., L. Childers and D. Margalit Cohomology of the hyperelliptic Torelli group

613

Childers, L. See T. Brendle Dijkstra, J. J. and K. I. S. Valkenburg Classifying Erdős type

spaces of higher descriptive complexity

725

Ding, L. and S. Gao On the Borelness of the intersection operation  783

Dolfi, S., R. Guralnick, C. E. Praeger and P. Spiga Coprime subdegrees for primitive permutation groups and completely reducible linear groups

745

Drensky, V. See S. Boumov Eisner, T. and T. Mátrai On typical properties of Hilbert space

operators

247

Fan, Q. Some C^{∗}-algebras properties preserved by tracial approximation

545

Fehm, A. and S. Petersen Hilbertianity of division fields of commutative algebraic groups

123

Feldheim, N. D. Zeroes of Gaussian analytic functions with translation-invariant distribution

317

Gao, S. See L. Ding

Gaudron, É. and G. Rémond Minima, pentes et algèbre tensorielle  565

Ghate, E. See D. Banerjee Guralnick, R. See S. Dolfi

Hazrat, R. The graded structure of Leavitt path algebras  833
Hazrat, R. and Z. Zhang Multiple commutator formulas  481

Hubard, I. See G. Araujo-Pardo Jiang, D., C. Nien and Y. Qin Generalized Shalika models of p-adic

SO_{4n} and functoriality

135

Kifer, Y. Nonconventional Poisson limit theorems  373

Kim, S. and S. T. Lee Pieri algebras for the orthogonal and symplectic groups

215

Kubiś, W. and S. Solecki A proof of uniqueness of the Gurariĭ space  449

Lee, S. J. See A. Barvinok Lee, S. T. See S. Kim Lefèvre, P., D. Li, H. Queffélec and L. Rodríguez-Piazza

Some new properties of composition operators associated with lens maps

801

Li, D. See P. Lefèvre López, F. J. See A. Alarcón Marchiafava, S. and R. Pantilie Twistor Theory for co-CR

quaternionic manifolds and related structures

347

Margalit, D. See T. Brendle Mátrai, T. See T. Eisner

Mozzochi, C. J. Linking numbers of modular geodesics  71
Murai, S. Face vectors of simplicial cell decompositions of manifolds  187

Nguyen, H. N. Quasisimple classical groups and their complex group algebras

973

Nien, C. See D. Jiang Novik, I. See A. Barvinok Oliveros, D. See G. Araujo-Pardo Pacini, M. and D. Testa Recovering plane curves of low degree from

their inflection lines and inflection points

283

Palka, K. Classification of singular ℚ-homology planes. I. Structure and singularities

37

Pantilie, R. See S. Marchiafava Petersen, S. See A. Fehm Plebanek, G. See A. Avilés Praeger, C. E. See S. Dolfi Qin, Y. See D. Jiang Qu, H. Finite non-elementary abelian p-groups whose number of sub-

groups is maximal

773

Queffélec, H. See P. Lefèvre Randriambololona, H. (2, 1)-separating systems beyond the prob-

abilistic bound

171

Regev, A. Lie superalgebras and some characters of S_{n}  31
Regev, O. Entropy-based bounds on dimension reduction in L_{1}  825

Rémond, G. See É. Gaudron Rodríguez, J. See A. Avilés Rodríguez-Piazza, L. See P. Lefèvre Saroglou, Ch. Minimal surface area position of a convex body is not

always an M-position

631

Schulte, E. See G. Araujo-Pardo Siciliano, S. and H. Usefi Lie identities on symmetric elements of

restricted enveloping algebras

999

Silberman, L. See N. Anantharaman Solecki, S. See W. Kubiś Spiga, P. See S. Dolfi Szpruch, D. Some irreducibility theorems of parabolic induction on

the metaplectic group via the Langlands–Shahidi method

897

Templier, N. and J. Tsimerman Non-split sums of coefficients of GL(2)-automorphic forms

677

Testa, D. See M. Pacini Tsimerman, J. See N. Templier Usefi, H. See S. Siciliano Valkenburg, K. I. S. See J. J. Dijkstra Zhang, Z. See R. Hazrat